OpenStudy (anonymous):

Please Help!! Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (anonymous):

@zzr0ck3r @Venny

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

So how would I wright that with the numbers? @jdoe0001

OpenStudy (jdoe0001):

find the dot product... you should have covered that and then find the magnitude

OpenStudy (jdoe0001):

check your material on the dot product of two vectors and their magnitude, it should be there

OpenStudy (anonymous):

(-5-4)*(-4-3) is the dot product?

OpenStudy (jdoe0001):

\(\large <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\)

OpenStudy (anonymous):

so (-5*-4)+(-4*-3)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

how do i find the magnitude?

OpenStudy (jdoe0001):

\(\large ||<a,b>||\implies \sqrt{a^2+b^2}\)

OpenStudy (anonymous):

(6.4)=u (5.3)=v??

OpenStudy (anonymous):

I'm not understanding this @jdoe0001

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

\(\bf u = <-5, -4>, v = <-4, -3> \\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right) \\ \quad \\ \theta = cos^{-1}\left(\cfrac{(-5\cdot -4)+(-4\cdot -3)}{\sqrt{(-5)^2+(-4)^2}\cdot \sqrt{(-4)^2+(-3)^2}}\right)\)

OpenStudy (jdoe0001):

that will give you a value, that will be \(\bf -1 \ge\ or \le 1\)

OpenStudy (anonymous):

ok thank you!! i got the answer!!

OpenStudy (jdoe0001):

yw